OpenStudy (anonymous):

A vessel has 6 kg of Hydrogen at pressure 'p' and temp '500K'. A small hole is made in it so that hydrogen leaks out . How much hydrogen leaks out if the final pressure is P/2 and temp falls to 300 K?

OpenStudy (experimentx):

what is the volume of 6 kg hydrogen?

OpenStudy (anonymous):

Not given in the question..

OpenStudy (experimentx):

this is not difficult. consider STP ... how do you find the volume of 6kg hydrogen?

OpenStudy (anonymous):

I am a vintage 1972 . It is my daughter who is posing the problem. I need to refer to text book to answer your question but it must something to do with M/m =n ( number of moles or volume into density =mass . pl help

OpenStudy (experimentx):

yeah ... 2g of hydrogen = 1M 6kg = 6000/2 1M = 22.4 L ----------------- using this relation you will find, 6kg = 3000 moles

OpenStudy (experimentx):

PV = 3000 R*500 P/2*V = n R*300 -------------- find 'n'

OpenStudy (anonymous):

PV/1500000R =PV/600 Rn or n = 1500000/600 =2500 or 2500 moles i.e 500 moles are leaked out. but options are limited to gms only . pl help

OpenStudy (experimentx):

2 g = 1 M

OpenStudy (anonymous):

1 moles = 2 gm 500 moles = 1000 gms=1 kg but options are 2g, 3g,5g &1g. Pl help

OpenStudy (anonymous):

as pv=nrt and v is constant or p/nrt is constant p1/n1rt1=p2/n2rt2 so,p/n1*500=p/2*n1*300>>>>as r is constant and other are given so, 600n2=500n1 n2=(5/6)n1 as initially it is 6kg i.e. 6000 moles so final is 5000moles i.e 5Kg :)

OpenStudy (anonymous):

so answer is 5:)

OpenStudy (anonymous):

But answers are (options) 2g, 3g, 5g & 1 g.

OpenStudy (experimentx):

answer can't be in grams!! grams is too low.

OpenStudy (anonymous):

yup i agree

OpenStudy (anonymous):

I think your original mass is supposed to be in grams in which case I believe the answer is 1 gram leaks out.

OpenStudy (anonymous):

yup i agree with u also

OpenStudy (experimentx):

using your data i got 1 kg ... if it were 6g, probably it would have been 1g

OpenStudy (anonymous):

Yeah my mistake . It is 6g and 6 kg as I mis-typed. Thanks a lot all of you.It good night at this end . Today in fact we celebrated Bijoya Dasami ( Last day of Durga Pooja ( for Bengali) and Dussera ( win over evil as celebrated in other parts of our country) So happy Dussera/ Subha Bijoya.

OpenStudy (anonymous):

I am back . I have got the answer as 1g leaked out.. Thanks all of you once again.